# CodeEnhance: A Codebook-Driven Approach for Low-Light Image Enhancement

### v1

@@ IV-C Comparison with State-of-the-Art Methods
As shown in Tabel. I, we compare the proposed CodeEnhance against SOTA LLIE methods regarding PSNR, SSIM, MAE, and LPIPS on the LOL, LSRW (Huawei and Nikon), FiveK, and SynLL datasets. In the case where the LL images contain much noise [1][24], some methods typically over-smooth the noise, leading to higher PSNR [2][4][3]. However, as shown in Figs. 6-9, they often lack high-frequency detail information and have poor perceptual quality, while the proposed CodeEnhance can obtain perceptually more convincing results, as indicated by the lowest LPIPS score. Figs. 6-9 show visual results comparisons. In a fully dark scene (Fig. 6 and Fig. 9), SNR and SMG improve overall brightness but introduce noise and artifacts, which damage the image quality. In scenes with uneven illumination (Fig. 7 and Fig. 8), methods like JED, KinD, and PairLIE enhance brightness better but suffer from color distortion and lack of details. SNR and SMG are susceptible to issues such as local color deviation and loss of texture. In contrast, our method enhances image illumination, not only suppressing noise and artifacts, but also accurately restoring color and texture information. Intuitively, the text in the figures appears significantly clearer, significantly improving the visual perception. This is accomplished by utilizing HQ image priors and incorporating advanced components such as SEM, CS, and IFT in our design.
Additionally, our method also has a superior performance in real scenes. Fig. 10 illustrates that our method achieves smooth processing in areas with a wide dynamic range of illumination changes (e.g., gradient sky colors and the transition between buildings and the sky, as depicted in Fig. 10. It effectively avoids artifacts and color deviation, resulting in visually pleasing and artifact-free enhancements. This demonstrates the robustness and effectiveness of our method in handling challenging scenarios with varying illumination conditions. Meanwhile, our method can adaptively enhance low-illumination areas instead of uniformly enhancing all dark areas. This adaptive approach helps to avoid amplifying noise and enables us to achieve excellent contrast while enhancing brightness. In addition, Table II demonstrates that our approach well trade-offs efficiency and image enhancement quality.
TABLE II: Runtime on LSRW Nikon dataset [38].
Methods JED [43] RetinexNet [1] KinD [44] Zero [45] EnGAN [12]
@@
Methods SNR [2] MIRNet v2 [4] PairLIE [46] SMG [3] Ours
@@ IV-D Low-light Object Detection
To assess the impact of our method in enhancing the performance of high-level tasks under low-light conditions, we utilize YOLOv7 [50] as the object detector and train it on the ExDark dataset, which was preprocessed using the LLIE methods. The comparison of Average Precision (AP) is presented in Table III, revealing that our method yields images that better facilitate the learning of the target detection algorithm. Furthermore, Fig. 11 demonstrates the effectiveness of our method in suppressing noise while increasing overall brightness in images with complex scenes. As a result, high-quality enhanced images obtained by the proposed CodeEnhance enable YOLOv7 to extract more valuable information and achieve better detection performance.
Figure 12: Visualized ablation of IFT on FiveK.
TABLE III: The mean Average Precision (mAP) comparisons of low-light object detection on ExDark [5].
LIME[42] 2016 82.75 77.59 73.92 90.12 83.6 67.22 69.51 74.01 78.29 64.05 80.95 54.24 74.68
JED[43] 2018 82.74 76.74 74.32 85.57 81.07 67.89 68.56 73.64 79.79 65.84 80.4 54.21 74.23
RetinexNet[1] 2018 81.97 77.59 74.08 85.68 79.61 54.92 63.32 71.28 66.39 59.08 77.59 51.17 70.22
KinD[44] 2019 81.13 78.04 74.97 87.96 82.69 59.95 66.62 72.4 72.25 63.69 80.28 50.44 72.54
Zero[45] 2020 83.44 77.22 75.17 89.7 82.4 62.67 66.8 73.17 75.68 64.16 80.68 50.22 73.44
EnGAN[12] 2021 81.68 76.39 74.83 89.44 80.16 63.81 65.57 72.53 77.18 64.24 80.55 51.51 73.16
SNR[2] 2022 83.65 77.44 72.97 89.01 82.7 62.64 66.51 71.02 74.91 64.74 80.19 50.55 73.03
MIRNet v2[4] 2023 85.02 79.35 74.27 90.68 82.91 66.36 67.87 70.14 79 64.76 80.67 55.6 74.72
PairLIE[46] 2023 82.87 77.38 72 88.11 82.38 62.03 70.23 70.44 76.31 64.1 80.11 54.95 73.41
SMG[3] 2023 82.05 76.51 71.71 88.07 79.11 64.97 66.84 68.32 77.69 63.16 77.88 53.78 72.51
@@ IV-E Ablation Study
Study of SEM. According to the Table IV, we first assess the proposed SEM. The results show that SEM effectively enhances the model’s performance. For example, when compared to Exp. (a) using the baseline model, PSNR and SSIM increased by 0.33 and 0.0152, respectively. These results indicate that SEM can acquire high-quality image features, thereby enhancing the accuracy of feature matching so as to improve overall model performance.
Study of IFT. IFT consists of a TFT and a CPT, and we will evaluate their effectiveness individually. Figs. 5, 12, and 13, and Table IV demonstrate that the proposed IFT and CPT can improve the performance of our CodeEnhance significantly. For instance, when compared to Exp. (b) in Table IV, there is a PSNR increase of 1.36 and an SSIM increase of 0.0167. These findings validate the effectiveness of our IFT in effectively controlling texture, color, and illumination information, thereby enhancing the overall model performance.
Study of CS. CS is designed to rectify distribution discrepancies among various datasets. As demonstrated in Table IV (Exp. (d) and (e)) and Table VI, CS has a beneficial impact on the model, resulting in an improvement of 0.16 in PSNR. This indicates that CS can effectively fine-tune the Codebook, adapting it to the current dataset and enhancing the model’s performance.
TABLE IV: Ablation studies of the CodeEnhance on FiveK [41] dataset. The Baseline model is built by a VQ-GAN [18]
@@
(e) 24.69 0.9023
TABLE V: Ablation studies of IFT, reference image, and codebook.
@@
PSNR 22.84 24.01 24.23 21.46 24.69
SSIM 0.8693 0.8893 0.8960 0.8227 0.9023
TABLE VI: Ablation studies of the $\lambda_{1}$ for $\mathcal{L}_{reg}$.
@@
PSNR 24.46 24.65 24.48 24.69
SSIM 0.9008 0.9006 0.9014 0.9023
@@ V Conclusion
In conclusion, we present a novel LLIE method, CodeEnhance, to obtain HQ images from LL images. We redefine LLIE as learning an image-to-code mapping between LL images and discrete codebook priors. Our method includes the Semantic Embedding Module (SEM) and Codebook Shift (CS) mechanism to enhance mapping learning by integrating semantic information and ensuring distribution consistency in feature matching. Additionally, the Interactive Feature Transformation (IFT) module improves texture and color information in image reconstruction, allowing interactive adjustment based on user preferences. Experimental results on real-world and synthetic benchmarks demonstrate that our approach improves LLIE performance in terms of quality, fidelity, and robustness to various degradations.

### v3

@@ IV-C Comparison with State-of-the-Art Methods
As shown in Tabel. I and Fig. 6, we compare the proposed CodeEnhance against SOTA LLIE methods regarding PSNR, SSIM, MAE, and LPIPS on the LOL, LSRW (Huawei and Nikon), FiveK, and SynLL datasets. In the case where the LL images contain much noise [11][3], some methods typically over-smooth the noise, leading to higher PSNR [4][5][6]. However, as shown in Figs. 7-10, they often lack high-frequency detail information and have poor perceptual quality, while the proposed CodeEnhance can obtain perceptually more convincing results, as indicated by the lowest LPIPS score. In a fully dark scene (Fig. 7 and Fig. 10), SNR and SMG improve overall brightness but introduce noise and artifacts, which damage the image quality. In scenes with uneven illumination (Fig. 8 and Fig. 9), methods like JED, KinD, and PairLIE enhance brightness better but suffer from color distortion and lack of details. SNR and SMG are susceptible to issues such as local color deviation and loss of texture. In contrast, our method enhances image illumination, not only suppressing noise and artifacts, but also accurately restoring color and texture information. Intuitively, the text in the figures appears significantly clearer, significantly improving the visual perception. This is accomplished by utilizing HQ image priors and incorporating advanced components such as SEM, CS, and IFT in our design.
Figure 12: Visual results on the LIME [54], MEF [10], and NPE [55] unpaired real-world dataset. Our method enhances the light of input images and maintains natural colors.
Figure 13: Visual results on the real scene. The results demonstrate the superior performance of our proposed CodeEnhance. Particularly, it excels in preserving texture details and effectively handling gradient change regions.
Figure 14: Local first-order differential values correspond to the first row of Fig. 13. The gradients of our results are continuous, broad, and smooth at the edges, indicating superior texture preservation and fewer artifacts
TABLE III: Runtime on LSRW Nikon dataset [47].
Methods JED RetinexNet KinD Zero EnGAN
@@
Methods SNR MIRNet v2 PairLIE SMG Ours
@@
To evaluate the generalization performance of the proposed CodeEnhance, we conduct experiments in self-captured night images and unpaired datasets, e.g., LIME [54], MEF [10], and NPE [55]. The results are presented in Table II, Fig. 11, Fig. 12, Fig. 13, and Fig. 14. Compared to other methods, CodeEnhance achieves superior generalization to unseen scenes. Fig. 12 shows that CodeEnhance effectively avoids artifacts and color deviation, resulting in visually pleasing and artifact-free enhancements. Fig. 13 and Fig. 14 illustrate that our method achieves smooth processing in areas with a wide dynamic range of illumination changes (e.g., gradient sky colors and the transition between buildings and the sky. This demonstrates the robustness and effectiveness of our method in handling challenging scenarios with varying illumination conditions. In addition, Table III demonstrates that our approach well trade-offs efficiency and image enhancement quality.
In summary, our improved results stem from the high-quality image priors and three key modules in CodeEnhance. First, the SEM fuses semantic information with low-level features in the encoder, effectively bridging the semantic gap between encoder outputs and the codebook for robust feature matching. Second, the CS mechanism adapts a pre-learned codebook to the unique characteristics of our LLIE dataset, ensuring consistency of distribution and highlighting the most relevant priors. Finally, the IFT module refines texture, color, and brightness by combining low-level encoder features with color guidance from reference images. Together, these design choices reduce uncertainty in the enhancement process and yield high-fidelity, perceptually convincing results.
TABLE IV: The mean Average Precision (mAP) comparisons of low-light object detection on ExDark [1].
LIME 82.75 77.59 73.92 90.12 83.6 67.22 69.51 74.01 78.29 64.05 80.95 54.24 74.68
JED 82.74 76.74 74.32 85.57 81.07 67.89 68.56 73.64 79.79 65.84 80.4 54.21 74.23
RetinexNet 81.97 77.59 74.08 85.68 79.61 54.92 63.32 71.28 66.39 59.08 77.59 51.17 70.22
KinD 81.13 78.04 74.97 87.96 82.69 59.95 66.62 72.4 72.25 63.69 80.28 50.44 72.54
Zero 83.44 77.22 75.17 89.7 82.4 62.67 66.8 73.17 75.68 64.16 80.68 50.22 73.44
EnGAN 81.68 76.39 74.83 89.44 80.16 63.81 65.57 72.53 77.18 64.24 80.55 51.51 73.16
SNR 83.65 77.44 72.97 89.01 82.7 62.64 66.51 71.02 74.91 64.74 80.19 50.55 73.03
MIRNet v2 85.02 79.35 74.27 90.68 82.91 66.36 67.87 70.14 79 64.76 80.67 55.6 74.72
PairLIE 82.87 77.38 72 88.11 82.38 62.03 70.23 70.44 76.31 64.1 80.11 54.95 73.41
SMG 82.05 76.51 71.71 88.07 79.11 64.97 66.84 68.32 77.69 63.16 77.88 53.78 72.51
@@ IV-D Low-light Object Detection
To assess the impact of our method in enhancing the performance of high-level tasks under low-light conditions, we utilize YOLOv7 [63] as the object detector and train it on the ExDark dataset, which was preprocessed using the LLIE methods. The comparison of Average Precision (AP) is presented in Table IV and Fig. 11(d), revealing that our method yields images that better facilitate the learning of the target detection algorithm. Furthermore, Fig. 15 demonstrates the effectiveness of our method in suppressing noise while increasing overall brightness in images with complex scenes. As a result, high-quality enhanced images obtained by the proposed CodeEnhance enable YOLOv7 to extract more valuable information and achieve better detection performance.
@@ IV-E Ablation Study
Study of SEM. According to the Table V, we first assess the proposed SEM. The results show that SEM effectively enhances the model’s performance. For example, when compared to Exp. (a) using the baseline model, PSNR and SSIM increased by 0.33 and 0.0152, respectively. These results indicate that SEM can acquire high-quality image features, thereby enhancing feature matching accuracy to improve overall model performance.
Study of IFT. IFT consists of a TFT and a CPT, and we will individually evaluate their effectiveness. Figs. 5 and 17, Table V and Table VI demonstrate that the proposed IFT and CPT significantly improve the performance of our CodeEnhance. For instance, when compared to Exp. (b) in Table V, there is an increase in PSNR of 1.36 and an increase in SSIM of 0.0167. These findings validate the effectiveness of our IFT in effectively controlling texture, color, and illumination information, thereby enhancing the overall model performance.
Figure 16: Visualized ablation of SEM on FiveK [53].
Figure 17: Visualized ablation of IFT (top) on FiveK [53]. and corresponding visual of feature maps (bottom). The higher the feature activation, the greater the feature score. It is evident that our IFT enhances CodeEnhance’s ability to capture more texture information.
Figure 19: Top 10 code activation frequencies from feature quantization on the FiveK dataset [53], measured under three conditions: VQ-GAN ground truth reconstruction (GT), LLIE by our model without Codebook Shift (w/o CS), and LLIE by our model with Codebook Shift (w/ CS).
TABLE V: Ablation studies of the CodeEnhance on FiveK [53] dataset. The Baseline model is built by a VQ-GAN [16]
@@
(e) 24.81 0.9086
TABLE VI: Ablation studies of IFT, reference image, and codebook on FiveK [53].
@@
PSNR 22.84 24.01 24.23 21.46 24.81
SSIM 0.8693 0.8893 0.8960 0.8227 0.9086
TABLE VII: Ablation studies of the $\lambda_{1}$ for $\mathcal{L}_{reg}$.
@@
PSNR 24.46 24.65 24.48 24.81
SSIM 0.9008 0.9006 0.9014 0.9086
Study of CS. CS is designed to rectify distribution discrepancies among various datasets. As demonstrated in Table V (Exp. (d) and (e)) and Table VII, CS has a beneficial impact on the model, resulting in an improvement of 0.16 in PSNR. Additionally, Fig. 19 illustrates the Top 10 code activation frequencies for the FiveK dataset. To measure how closely the sampling trends align with the ground truth, we use the Pearson correlation coefficient. Our results show that the model with Codebook Shift (w/ CS, p = 0.9505) exhibits a stronger alignment with the ground truth compared to the model without Codebook Shift (w/o CS, p = 0.8791), indicating that Codebook Shift effectively guides the sampling process toward the ground-truth distribution.
Study of Codebook and Reference Image. Table VI presents an ablation study evaluating the contributions of both the reference image and the codebook. The reference image provides brightness and contrast cues that guide the enhancement process, while the codebook reframes the task as “image-to-code” and leverages high-quality priors to reduce uncertainty in the process. As shown, removing the reference image (w/o Refer Img) still produces competitive results but remains slightly below the final version, highlighting the benefit of including external brightness/contrast information. Meanwhile, omitting the codebook (w/o codebook) causes a notable performance drop, confirming that high-quality priors are critical for preserving reconstruction fidelity.
@@ V Conclusion
In conclusion, we present a novel LLIE method, CodeEnhance, to obtain HQ images from LL images. We redefine LLIE as learning an image-to-code mapping between LL images and discrete codebook priors. Our method includes the Semantic Embedding Module (SEM) and Codebook Shift (CS) mechanism to enhance mapping learning by integrating semantic information and ensuring distribution consistency in feature matching. Additionally, the Interactive Feature Transformation (IFT) module improves texture and color information in image reconstruction, allowing interactive adjustment based on user preferences. Experimental results on real-world and synthetic benchmarks demonstrate that our approach improves LLIE performance in terms of quality, fidelity, and robustness to various degradations. However, our method is built on a GAN framework, whose image generation capability is not as strong as diffusion models. To further improve LLIE, we plan to integrate vector quantization with a diffusion model. The high-quality image representations learned by vector quantization can serve as effective prompts for the diffusion model, potentially enabling even more advanced enhancement capabilities.
Acknowledgement This work was supported in part by the Natural Science Foundation of China (No. 61976145, No. 62076164, and No. 62272319), Guangdong Basic and Applied Basic Research Foundation (No. 2023A1515010677 and No. 2021A1515011861), and Shenzhen Science and Technology Program (No. JCYJ20210324094413037)